OpenStudy (anonymous):

The area of circle is halved when its radious is decreased by n. What can be the value of radious of the circle

OpenStudy (anonymous):

The formula for the area of a circle is: \(A = \pi r^2\) Let's say the radius here is 10. \(A = \pi (10^2)\) \(A = 100\pi\) Half of this would be \(50 \pi\). Let's plug this into the formula for the area: \(A = \pi r^2\) \(50\pi = \pi r^2\) \(50 = r^2\) \(r = \sqrt{50}\) \(r \approx 7.07\) Now we subtract 10 - 7.07 What do you get?